OpenStudy (anonymous):

how many 5/8 are in 21 7/8

OpenStudy (anonymous):

21 7/8 divided by 5/8 is 35

OpenStudy (anonymous):

start with something simple: how many 1/4 are in one =4 how many 1/3 are in one=3 how did you get that? 1 divided by 1/4 = 1 x 4/1 = 4 similarly 1 divided by 1/3 = 1 x 3/1 = 3 so, 21 and 7/8 = (21x8) + (7/8) = (168/8) +(7/8) =175/8 so, 175/8 divided by 5/8 = 175/8 x (8/5) you will notice the 8's cancel an you are left ith 175/5 =35